# Beyond Seen Data: Improving KBQA Generalization Through Schema-Guided Logical Form Generation

### v2

@@ Abstract
Knowledge base question answering (KBQA) aims to answer user questions in natural language using rich human knowledge stored in large KBs. As current KBQA methods struggle with unseen knowledge base elements at test time, we introduce SG-KBQA: a novel model that injects schema contexts into entity retrieval and logical form generation to tackle this issue. It uses the richer semantics and awareness of the knowledge base structure provided by schema contexts to enhance generalizability. We show that SG-KBQA achieves strong generalizability, outperforming state-of-the-art models on two commonly used benchmark datasets across a variety of test settings. Our source code is available at https://github.com/gaosx2000/SG_KBQA.
@@
Shengxiang Gao           Jey Han Lau           Jianzhong Qi School of Computing and Information Systems, The University of Melbourne shengxiang.gao1@student.unimelb.edu.au {jeyhan.lau, jianzhong.qi}@unimelb.edu.au
@@ 1 Introduction
Knowledge base question answering (KBQA) aims to answer user questions expressed in natural language with information from a knowledge base (KB). This offers user-friendly access to rich human knowledge stored in large KBs such as Freebase Bollacker et al. (2008), DBPedia Auer et al. (2007) and Wikidata Vrandečić and Krötzsch (2014), and it has broad applications in QA systems Zhou et al. (2018), recommender systems Guo et al. (2022), and information retrieval systems Jalota et al. (2021).
Figure 1: Example of KBQA and SP-based solutions.
State-of-the-art (SOTA) solutions often take a semantic parsing (SP)-based approach. They translate an input natural language question into a structured, executable form (AKA logical form Lan et al. (2021)), which is then executed to retrieve the question answer. Figure 1 shows an example. The input question, Who is the author of Harry Potter, is expressed using the S-expression Gu et al. (2021) (a type of logical form), which is formed by a set of functions (e.g., JOIN) operated over elements of the target KB (e.g., entity m.078ffw refers to book series Harry Potter, book.author a class of entities, and book.literary_series.author a relation in Freebase).
A key challenge here is to learn a mapping between mentions of entities and relations in the input question to corresponding KB elements to form the logical form. Meanwhile, the mapping of KB element compositions has to adhere to the structural constraints (schema) of the KB. The schema defines entities’ classes and the relationships between these classes within the KB. Take the KB subgraph in Figure 1 as an example, the relationship between the entity Harry Potter and the entity J.K. Rolwing is defined by the relation book.literary_series.author between their respective classes (i.e., class book.literary_series and class book.author).
Figure 2: Relation-guided entity mention detection and schema-guided logical form generation.
@@
However, due to the vast number of entities, relations, classes, and their compositions, it is difficult (if not impossible) to train a model with all feasible compositions of the KB elements. For example, Freebase Bollacker et al. (2008) has over 39 million entities, 8,000 relations, and 4,000 classes. Furthermore, some KBs (e.g., NED Mitchell et al. (2018)) are not static as they continue to grow.
@@
A few studies consider model generalizability to non-I.I.D. settings, where the test set contains schema items (i.e., relations and classes) or compositions that are unseen during training (i.e., zero-shot and compositional generalization, respectively). In terms of methodology, these studies typically use ranking-based or generation-based models. Ranking-based models Gu et al. (2021, 2023) retrieve entities relevant to the input question and then, starting from them, perform path traversal in the KB to obtain the target logical form by ranking. Generation-based models Shu et al. (2022); Zhang et al. (2023) retrieve relevant KB contexts (e.g., entities and relations) for the input question, and then feed these contexts into a Seq2Seq model together with the input question to generate the logical form.
We observe that both types of models terminate their entity retrieval prematurely, such that each entity mention in the input question is mapped to only a single entity before the logical form generation stage. As a result, the logical form generation stage loses the freedom to explore the full combination space of relations and entities. This leads to inaccurate logical forms (as validated in our study).
To address this issue, our strategy is to defer entity disambiguation — i.e., to determine the most relevant entity for an entity mention (Section 2) — to the logical form generation stage. This allows our model to explore a larger combination space of the relations and entities, and ultimately leads to stronger model generalizability because low-ranked (but correct) relations or entities would still be considered during generation. We call our approach SG-KBQA (schema-guided logical form generator for KBQA). Concretely, SG-KBQA follows the generation-based approach but with deferred entity disambiguation. As shown in Figure 2, it feeds the input question, the retrieved candidate relations and entities, plus their corresponding schema information (the domain and range of classes of relations and entities; Section 4) into a large language model (LLM) for logical form generation. The schema information reveals the connectivity between the candidate relations and entities, hence guiding the LLM to uncover their correct combination in the large search space.
Further exploiting the schema-guided idea, we propose a relation-guided module for SG-KBQA  to enhance its entity mention detection from the input question. As shown in Figure 2, this module adapts a Seq2Seq model to generate logical form sketches based on the input question and candidate relations, where relations, classes, and literals are masked by special tokens, such that the entity mentions can be identified more easily without confusions caused by these elements.
To summarise:
  • We introduce SG-KBQA to solve the KBQA problem under non-I.I.D. settings, where test input contains unseen schema items or compositions during training.
  • We propose to defer entity disambiguation to logical form generation, and additionally guide this generation step with corresponding schema information, allowing us to explore a larger combination space of relations and entities to consider unseen relations, entities, and compositions. We further propose a relation-guided module to strengthen entity retrieval by generating logical form sketches.
@@
  • We conduct experiments on two popular benchmark datasets and find SG-KBQA outperforming SOTA models on both datasets. In particular, on non-I.I.D GrailQA our model tops all three leaderboards for the overall, zero-shot, and compositional generalization settings, outperforming SOTA models by 3.3%, 2.9%, and 4.0% (F1) respectively.
@@ 2 Related Work
Knowledge Base Question Answering
Most KBQA solutions use information retrieval-based (IR-based) or semantic parsing-based (SP-based) methods Wu et al. (2019); Lan et al. (2021). IR-based methods construct a question-specific subgraph starting from the retrieved entities (i.e., the topic entities). They then reason over the subgraph to derive the answer. SP-based methods focus on transforming input questions into logical forms, which are then executed to retrieve answers. SOTA solutions are mostly SP-based, as detailed next.
KBQA under I.I.D. Settings
Recent KBQA studies under I.I.D. settings fine-tune LLMs to map input questions to rough KB elements and generate approximate logical form drafts Luo et al. (2024); Wang and Qin (2024). The approximate (i.e., inaccurate or ambiguous) KB elements are then aligned to exact KB elements through a subsequent retrieval stage. These solutions often fail over test questions that refer to KB elements unseen during training. While we also use LLMs for logical form generation, we ground the generation with retrieved relations, entities, and schema contexts, thus addressing the non-I.I.D. issue.
@@
Ranking-based methods start from retrieved entities, traverse the KB, and construct the target logical form by ranking the traversed paths. Gu et al. (2021) enumerate and rank all possible logical forms within two hops of retrieved entities, while Gu et al. (2023) incrementally expand and rank paths from retrieved entities.
Generation-based methods transform an input question into a logical form using a Seq2Seq model (e.g., T5 Raffel et al. (2020)). They often use additional contexts beyond the question to augment the input of the Seq2Seq model and enhance its generalizability. For example, Ye et al. (2022) use top-5 candidate logical forms enumerated from retrieved entities as the additional context. Shu et al. (2022) further use top-ranked relations, disambiguated entities, and classes (retrieved separately) as the additional context. Zhang et al. (2023) use connected pairs of retrieved KB elements.
Our SG-KBQA is generation-based. We use schema contexts (relations and classes) from retrieved relations and entities, rather than separate class retrieval (as in Shu et al. (2022)) which could introduce noise. We also defer entity disambiguation to the logical form generation stage, thus avoiding error propagation induced by premature entity disambiguation without considering the generation context, as done in existing works outlined below.
@@
KBQA entity retrieval typically has three steps: entity mention detection, candidate entity retrieval, and entity disambiguation. BERT Devlin et al. (2019)-based named entity recognition is widely used for entity mention detection from input questions. To retrieve KB entities corresponding to entity mentions, the FACC1 dataset Gabrilovich et al. (2013) is commonly used, which contains over 10 billion surface forms (with popularity scores) of Freebase entities. Gu et al. (2021) use the popularity scores for entity disambiguation, while Ye et al. (2022) and Shu et al. (2022) adopt a BERT reranker.
@@ 3 Preliminaries
Figure 3: Overview of SG-KBQA. The model has two stages: retrieval and generation. In the retrieval stage, we first retrieve and rank candidate relations based on the input question $q$ (①). Using $q$ and the top-ranked candidate relations $R_{q}$, we generate logical form sketches and extract entity mentions from them (②). Based on the entity mentions and retrieved relations, we retrieve candidate entities from the KB (③) and rank them (the top-ones being $E_{q}$, ④). In the generation stage, $q$, $R_{q}$, $E_{q}$, and their class contexts, are fed into a fine-tuned language model for logical form generation (⑤). Here, the colored modules come with our new design.
A graph structured-KB $\mathcal{G}$ is composed of a set of relational facts $\{\langle s,r,o\rangle|s\in\mathcal{E},r\in\mathcal{R},o\in\mathcal{E}\cup\mathcal{L}\}$ and an ontology $\{\langle c_{d},r,c_{r}\rangle|c_{d},c_{r}\in\mathcal{C},r\in\mathcal{R}\}$. Here, $\mathcal{E}$ denotes a set of entities, $\mathcal{R}$ denotes a set of relations, and $\mathcal{L}$ denotes a set of literals, e.g., textual labels, numerical values, or date-time stamps. In a relational fact $\langle s,r,o\rangle$, $s\in\mathcal{E}$ is the subject, $o\in\mathcal{E}\cup\mathcal{L}$ is the object, and $r\in\mathcal{R}$ represents the relationship between the subject and the object.
The ontology defines the rules governing the composition of relational facts within $\mathcal{G}$. In its formulation, $\mathcal{C}$ denotes a set of classes, each of which defines a set of entities (or literals) sharing common properties (relations). Note that an entity can belong to multiple classes. In an ontology triple $\langle c_{d},r,c_{r}\rangle$, $c_{d}$ is called a domain class, and it refers to the class of subject entities that satisfy relation $r$; $c_{r}$ is called the range class, and it refers to the class of object entities or literals satisfying $r$. Each ontology triple can be instantiated as a set of relational facts. In Figure 1, <book.literary_series, book.literary_series.author, book.author> is an ontology triple. An instance of it is <Harry Potter, book.literary_series.author, J.K. Rowling>, where Harry Potter is an entity that belongs to class book.literary_series.
@@
Given a KB $\mathcal{G}$ and a question $q$ expressed in natural language, i.e., a sequence of word tokens, knowledge base question answering (KBQA) aims to find a subset (the answer set) $\mathcal{A}\subseteq\mathcal{E}\cup\mathcal{L}$ of elements from $\mathcal{G}$ that — with optional application of some aggregation functions (e.g., count) — answers $q$.
@@
We solve the KBQA problem by translating the input question $q$ into a structured query that can be executed on $\mathcal{G}$ to fetch the answer set $\mathcal{A}$. Following previous works Shu et al. (2022); Ye et al. (2022); Gu et al. (2023); Zhang et al. (2023), we use logical form as the structured query language, expressed with the S-expression Gu et al. (2021). The S-expression offers a readable representation well-suited for KBQA. It uses set semantics where functions operate on entities or entity tuples without requiring variables Ye et al. (2022). Figure 1 shows an example: the S-expression of the given question Who is the author of Harry Potter? is (AND book.author (JOIN (R book.literary_series.author) m.078ffw)). This S-expression queries a set of entities that belong to the class book.author from the objects of triples whose subject entity is m.078ffw while the relation is book.literary_series.author. More details about the S-expression is in Appendix A.
@@ 4 The SG-KBQA Model
As shown in Figure 3, SG-KBQA follows the common structure of generation-based models. It has two overall stages: relation and entity retrieval and logical form generation. We propose novel designs in both stages to strengthen model generalizability.
In the relation and entity retrieval stage (Section 4.1), SG-KBQA retrieves candidate relations and entities from KB $\mathcal{G}$ which may be relevant to the input question $q$. It starts with a BERT-based relation ranking model to retrieve candidate relations relevant to $q$. Together with $q$, the set of top-ranked candidate relations are fed into a novel, relation-guided Seq2Seq model to generate logical form sketches that contain entity mentions while masking the relations and classes. We harvest the entity mentions and use them to retrieve candidate entities from $\mathcal{G}$. We propose a combined relation-based strategy to prune the entities (as there may be many). The remaining entities are ranked by a BERT-based model, indicating their likelihood of being the entity that matches each entity mention.
@@
Leveraging relations to guide both entity mention extraction and candidate entity pruning enhances the model generalizability over entities unseen during training. This in turn helps the logical form generation stage to filter false positive matches for unseen relations or their combinations.
@@
In the logical form generation stage (Section 4.2), SG-KBQA feeds $q$, the top-ranked relations and entities (corresponding to each entity mention), and the schema contexts (i.e., domain and range classes of the relations and classes of the entities), into an adapted LLM to generate the logical form and produce answer set $\mathcal{A}$.
Our schema-guided logical form generation procedure is novel in that it takes (1) multiple candidate entities (instead of one in existing models) for each entity mention and (2) the schema contexts as the input. Using multiple candidate entities essentially defers entity disambiguation, which is usually done in the retrieval stage by existing models Shu et al. (2022); Gu et al. (2023), to the generation stage, thus mitigating error propagation. This strategy also brings challenges, as the extra candidate entities (which are ambiguous as they often share the same name) may confuse the logical form generation model. We address the challenges with the schema contexts, which instruct the model the connectivity structures between the candidate entities and relations. The connectivity structures further help SG-KBQA generalize to unseen entities, relations, or their combinations.
4.1 Relation and Entity Retrieval
Relation Retrieval

### v3

@@ Abstract
Knowledge Base Question Answering (KBQA) aims to answer user questions in natural language using rich human knowledge stored in large KBs. As current KBQA methods struggle with unseen knowledge base elements and their novel compositions at test time, we introduce SG-KBQA — a novel model that injects schema contexts into entity retrieval and logical form generation to tackle this issue. It exploits information about the semantics and structure of the knowledge base provided by schema contexts to enhance generalizability. We show that SG-KBQA achieves strong generalizability, outperforming state-of-the-art models on three commonly used benchmark datasets across a variety of test settings. Our source code is available at https://github.com/gaosx2000/SG_KBQA.
@@
Shengxiang Gao      Jey Han Lau      Jianzhong Qithanks: Corresponding author. The University of Melbourne shengxiang.gao1@student.unimelb.edu.au {laujh, jianzhong.qi}@unimelb.edu.au
@@ 1 Introduction
Knowledge Base Question Answering (KBQA) aims to answer user questions expressed in natural language with information from a knowledge base (KB). This offers user-friendly access to rich human knowledge from large KBs such as Freebase (Bollacker et al., 2008), DBPedia (Auer et al., 2007) and Wikidata (Vrandečić and Krötzsch, 2014), and it has broad applications in QA (Zhou et al., 2018), recommendation (Guo et al., 2022), and information retrieval (Jalota et al., 2021).
Semantic Parsing (SP) has been shown to be an effective method for KBQA, where the core idea is to translate the input natural language question into a structured logical form (e.g., SPARQL or S-Expression (Gu et al., 2021)), which is then executed to yield the question answer.
Figure 1: Schema-guided entity retrieval (top) and logical form generation (bottom). Green arrows and boxes highlight schema-level connections (overlapping classes) among KB elements. “w schema” and “w/o schema” logical forms denote whether the composition of KB elements adheres to the KB schema, respectively.
A key challenge here is to learn a mapping between mentions of entities and relations in the input question to corresponding KB elements to form the logical form. Given a large number of entities and relations in ambiguous surface forms, and the flexibility in questions expressed in natural language, this mapping process typically yields a set of candidate entities (relations) for each mentions of entities (relations). The challenge then becomes to uncover the right composition of entities and relations from the sets of candidates.
Figure 1 shows an example. Two entities named Harry Potter (a book series and a main character in them) and two authorship relations of similar names were identified as candidates. Combining the top-ranked entity (the book series) with the top-ranked relation book.author.works_written (book authorship) yields an invalid and unexecutable logical form, as the entity is a book series, not a book.
@@
Due to the vast number of KB elements and their compositions, it is difficult (if not impossible) to train a model with all feasible compositions of KB elements that might be queried. For example, Freebase (Bollacker et al., 2008) has over 39 million entities, 8,000 relations, and 4,000 classes. Furthermore, some KBs (e.g., NED (Mitchell et al., 2018)) are not static as they continue to grow. When a KB element composition is unseen at training, errors like the example above may occur.
@@
A few studies consider model generalizability to non-I.I.D. settings, where the test set contains schema items (i.e., relations, classes and functions) or their compositions that are unseen during training (i.e., zero-shot and compositional generalization, respectively). They propose retrieval methods to retrieve KB elements or compositions more relevant to input questions, and use them to construct logical forms (Shu et al., 2022; Gu et al., 2023).
Despite these efforts, achieving compositional and zero-shot generalization remains challenging: (1) Entity retrieval is a bottleneck. Current entity retrieval methods often fail to accurately detect entities mentioned in questions containing schema items unseen at training. This is because such items introduce novel contextual patterns in questions, making it difficult to identify the correct boundary of entity mentions. The resulting entity retrieval errors propagate and lead to errors in the logical forms generated. (2) Schema-level connection between KB elements have been missed. Existing methods are not explicitly trained to capture compositions of KB elements that are feasible based on their schema. Instead, they tend to reproduce KB element compositions observed at training, making them difficult to generalize to unseen compositions.
To address these challenges, we propose a schema-guided model for KBQA (SG-KBQA), that incorporates KB schema to guide both entity retrieval and logical form generation. Unlike previous approaches that initiate the pipeline with entity retrieval, SG-KBQA adopts a schema-first principle, prioritizing schema understanding as the foundation for downstream logical form generation.
SG-KBQA begins with relation retrieval, employing a pre-trained language model (PLM)-based (Devlin et al., 2019) retriever to retrieve top-ranked relations from the KB that are most relevant to the input question. Benefiting from the generalization capability of pre-trained language models (and that there are much fewer relations and their surface form variants than entities in a KB), relations that are semantically similar to the question—yet unseen at training—can still be included among the top-ranked retrieved relations (as validated in our study). Then, as illustrated in Figure 1, we introduce a schema-guided entity retrieval (SER) module. This module employs a logical form sketch parser that converts the input question and retrieved relations into logical form sketches by a Seq2Seq model. These top-ranked relations provide schema context that is relevant to the question but not observed at training, hence helping the model distinguish actual entity mentions from unseen schema items in the question. More precise entity mentions are then extracted from the generated sketches, thereby improving the zero-shot generalizability of entity retrieval.
To further mitigate error propagation between the retrieval and generation stages, we defer entity disambiguation to the logical form generation stage. For each entity mention, all top-ranked matched candidates are retained as candidate entities for the preceding generation stage.
Further exploiting the schema-guided idea, we propose a schema-guided logical form generation (SLFG) module that fine-tunes a large language model (LLM) to reason over feasible compositions of KB elements based on their underlying schema contexts. As Figure 1 shows, we feed the input question, the retrieved candidate relations and entities, plus their corresponding schema contexts, i.e., the (domain and range) classes of the relations and entities, into the LLM for logical form generation. The domain and range classes of a relation refer to the classes to which its subject and object entities belong. Together with the class of the candidate entities, they provide explicit training signals to guide the LLM to look for KB elements that can be connected together (and hence are more likely to form executable logical forms). As a result, our SLFG module generalizes to compositions of KB elements unseen at training.
@@
To summarise: (1) We introduce SG-KBQA to solve the KBQA problem under non-I.I.D. settings, where test input contains unseen schema items or their compositions during training. (2) We introduce schema-guided modules for entity retrieval and logical form generation with deferring entity disambiguation to enhance both compositional and zero-shot generalization. These modules can also be incorporated into existing SP-based KBQA systems to improve their generalization performance. (3) We conduct experiments on three popular benchmark datasets and find SG-KBQA outperforming SOTA models. In particular, on non-I.I.D GrailQA our model tops all three leaderboards for the overall, zero-shot, and compositional generalization settings, outperforming SOTA models by 3.3%, 2.9%, and 4.0% (F1) respectively.
@@ 2 Related Work
Knowledge Graph Question Answering
Early KBQA solutions can be widely categorized as Information Retrieval-based (IR-based) (He et al., 2021; Zhang et al., 2022) or Semantic Parsing-based (SP-based) solutions Cao et al. (2022); Ye et al. (2022).
Benefiting from the strong natural language understanding and reasoning abilities of LLMs, recent LLM-based KBQA methods have achieved promising results. A branch of work employs LLMs to produce reasoning trajectories prior to producing final answers. For example, Luo et al. (2024b) and Luo et al. (2025) fine-tune LLMs to directly generate the reasoning trajectories in a single pass. In contrast, Sun et al. (2024), Liu et al. (2024) and Sui et al. (2025) follow an agentic framework, where the model performs step-by-step tool invocations to iteratively traverse the subgraph starting from topic entities, assuming that topic entities are given.
Others follow the SP-based paradigm, using LLMs to generate approximate logical form sketches through few-shot in-context learning or fine-tuning (Cao et al., 2022; Li et al., 2023, 2024; Luo et al., 2024a; Wang and Qin, 2024). The inaccurate or ambiguous KB elements in the generated sketches are further refined through a retrieval stage, aligning them with actual KB elements to construct complete logical forms.
However, these methods often fail to generalize over test questions that refer to KB elements or their compositions unseen during training, or when the topic entities are not known. Our SG-KBQA improves KBQA generalizability through a schema-guided approach. While we also use LLMs to generate logical form sketches, we incorporate retrieved relations to guide sketch generation for entity mention extraction, thereby improving the generalizability of entity retrieval, while we do not refine these sketches to produce the final output logical forms.
@@
Ranking-based methods Gu et al. (2021, 2023) start from retrieved entities, traverse the KB, and construct the target logical form by ranking the traversed paths.
Generation-based methods transform an input question into a logical form using a Seq2Seq model (e.g., T5 Raffel et al. (2020)). They often use additional contexts beyond the question to augment the input of the Seq2Seq model and enhance its generalizability. For example, Ye et al. (2022) use the top-5 candidate logical forms enumerated from the retrieved entities. Shu et al. (2022) further use top-ranked relations, disambiguated entities, and classes (retrieved separately). Zhang et al. (2023) use connected pairs of retrieved KB elements.
Our SG-KBQA adopts a generation-based approach, training the LLM to reason over candidate entities and relations, using their schema contexts (i.e., classes) to infer connectivity. This enables the model to compose novel logical forms without seeing them at training, hence generalizing better in larger, noisier search spaces. Additionally, we defer entity disambiguation to the generation stage, mitigating error propagation caused by early disambiguation without context.
@@
KBQA entity retrieval typically has three steps: entity mention detection, candidate entity retrieval, and entity disambiguation. BERT Devlin et al. (2019)-based named entity recognition is used for entity mention detection from input questions. To retrieve KB entities for the entity mentions, the FACC1 dataset Gabrilovich et al. (2013) is often used, with over 10 billion surface forms and popularity scores of Freebase entities. Gu et al. (2021) use popularity scores for entity disambiguation, while Ye et al. (2022) and Shu et al. (2022) adopt a BERT reranker.
@@ 3 Preliminaries
Figure 2: Overview of SG-KBQA. The model consists of two novel modules: schema-guided entity retrieval (SER) and schema-guided logical form generation (SLFG). Given a question $q$, the model first retrieves and ranks candidate relations (①). In SER, $q$ and the top-ranked relations $R_{q}$ are used to generate logical form sketches and extract entity mentions (②). Based on these mentions and $R_{q}$, the model retrieves and ranks candidate entities (③), producing the top entities $E_{q}$ (④). Entity disambiguation is deferred by directly passing $E_{q}$ to SLFG. In SLFG, $q$, $R_{q}$, $E_{q}$, and their class contexts are fed into a fine-tuned language model for logical form generation (⑤).
A graph structured-KB $\mathcal{G}$ is composed of a set of relational facts $\{\langle s,r,o\rangle|s\in\mathcal{E},r\in\mathcal{R},o\in\mathcal{E}\cup\mathcal{L}\}$ and an ontology $\{\langle c_{d},r,c_{r}\rangle|c_{d},c_{r}\in\mathcal{C},r\in\mathcal{R}\}$. Here, $\mathcal{E}$ denotes a set of entities, $\mathcal{R}$ a set of relations, and $\mathcal{L}$ a set of literals, e.g., textual labels or numerical values. In a relational fact $\langle s,r,o\rangle$, $s\in\mathcal{E}$ is the subject, $o\in\mathcal{E}\cup\mathcal{L}$ is the object, and $r\in\mathcal{R}$ represents the relation between the two.
The ontology defines the rules governing the composition of relational facts within $\mathcal{G}$: $\mathcal{C}$ denotes a set of classes, each of which defines a set of entities (or literals) sharing common properties (relations). Note that an entity can belong to multiple classes. In an ontology triple $\langle c_{d},r,c_{r}\rangle$, $c_{d}$ is the domain class, i.e., the class of subject entities that satisfy relation $r$; $c_{r}$ is the range class, i.e., the class of object entities or literals satisfying $r$. Each ontology triple can be instantiated a set of relational facts, with an example provided in Appendix A.
@@
Given a KB $\mathcal{G}$ and a question $q$ expressed in natural language, i.e., a sequence of word tokens, KBQA aims to find a subset (the answer set) $\mathcal{A}\subseteq\mathcal{E}\cup\mathcal{L}$ of elements from $\mathcal{G}$ that — with optional application of some aggregation functions (e.g., count) — answers $q$.
@@
We approach the KBQA problem by translating question $q$ into a structured query that can be executed on $\mathcal{G}$ to fetch the answer set $\mathcal{A}$. Following previous works Shu et al. (2022); Gu et al. (2023); Zhang et al. (2023), we use logical form as the structured query language, expressed in S-expression Gu et al. (2021). S-expression offers a readable representation well-suited for KBQA. It uses set semantics where functions operate on entities or entity tuples without requiring variables Ye et al. (2022), with more details in Appendix A.
@@ 4 The SG-KBQA Model
SG-KBQA takes a generation-based approach overall. It introduces two novel modules: Schema-guided Entity Retrieval (SER) and Schema-guided Logical Form Generation (SLFG), designed to enhance generalizability and shown in Figure 2.
SG-KBQA starts with relation retrieval, where a BERT-based relation ranking model retrieves candidate relations and entities from the KB $\mathcal{G}$ that are potentially relevant to the question $q$.
In SER, $q$ and the top-ranked candidate relations are passed into a logical form parser (i.e. a Seq2Seq model) to generate logical form sketches that contain entity mentions while masking out relations and classes. The retrieved relations provide the most relevant—and potentially unseen—relations as additional schema context, enabling the model to identify boundaries of entity mentions more accurately as explained in Section 1. We then harvest these entity mentions and use them to retrieve candidate entities from $\mathcal{G}$, thereby improving the non-I.I.D. generalizability of entity retrieval.
@@
To further improve the accuracy of entity retrieval, we propose a combined schema-based pruning strategy to filter out unlikely candidates, as a single mention may correspond to multiple entities. The remaining entities are then ranked by a BERT-based model, which estimates the likelihood of each entity being the correct match for a mention. Leveraging relations—a type of schema item—to guide both entity mention extraction and candidate entity pruning enhances model generalizability over entities unseen at training. This in turn helps logical form generation to filter false positive matches for unseen relations or their compositions.
@@
In SLFG, SG-KBQA feeds $q$, the top-ranked relations and entities (corresponding to each mention), and the schema contexts, i.e., their class information, into an adapted LLM to generate the logical form and produce answer set $\mathcal{A}$. SLFG is novel in that it takes (1) multiple candidate entities (instead of one in existing models) for each mention and (2) the schema contexts as the input.
By deferring entity disambiguation to the generation stage, our approach helps mitigate error propagation that often arises from early-stage disambiguation. This strategy also brings challenges, as the extra candidate entities (which often share the same or similar names) introduce noise in SLFG.
To address this and enhance generalizability to unseen compositions of KB elements, we incorporate schema context in SLFG. The LLM is fine-tuned to generate logical forms that consist of valid KB element compositions, as connected based on the (domain and range) classes of the relations and entities.
As a result, the model is able to select correct compositions from a noisy KB element space, even under non-I.I.D settings.
4.1 Relation Retrieval